An alloy of metals is 25% copper. Another alloy is 50% copper. How much of each alloy should be used to make 1000 grams of an alloy that is 45% copper?
Answer by stanbon(75887)

An alloy of metals is 25% copper. Another alloy is 50% copper. How much of each alloy should be used to make 1000 grams of an alloy that is 45% copper?
--------
Equation:
copper + copper = copper
0.25x + 0.50(1000-x) = 0.45*1000
----
Multiply thru by 100 to get:
25x + 50*1000 - 50x = 45*1000
-25x = -5*1000
x = 200 grams (amt. of 25% copper needed)
----
1000-x = 800 grams (amt. of 50% copper needed)
===============================
Cheers,
Stan H.